i dont know how to find the awnser to this
write the equation in standard form for the ellipse with foci (3,0) and (-3,0) and the endpoints on the vertical axis of (0,6) and (0,-6)
thanks i hope you can help!

Answer by stanbon(75887)

write the equation in standard form for the ellipse with foci (3,0) and (-3,0) and the endpoints on the vertical axis of (0,6) and (0,-6)
^Draw the picture.
The major axis is on the x-axis
Foci at (3,0) and (-3,0) means c=3
Endpts (0,6) and (0,-6) on the y-axis means b=6
For an ellipse a^2 = b^2 + c^2
a^2=6^2 + 3^2
a^2=45
EQUATION:
(x^2/45) + (y^2/36) =1
Cheers,
Stan H.